### v2

# MAPS: A Multilingual Benchmark for Agent Performance and Security

Abstract
Agentic AI systems, which build on Large Language Models (LLMs) and interact with tools and memory, have rapidly advanced in capability and scope. Yet, since LLMs have been shown to struggle in multilingual settings, typically resulting in lower performance and reduced safety, agentic systems risk inheriting these limitations. This raises concerns about the accessibility of such systems, as users interacting in languages other than English may encounter unreliable or security-critical agent behavior. Despite growing interest in evaluating agentic AI, existing benchmarks focus exclusively on English, leaving multilingual settings unexplored. To address this gap, we propose MAPS, a multilingual benchmark suite designed to evaluate agentic AI systems across diverse languages and tasks. MAPS builds on four widely used agentic benchmarks — GAIA (real-world tasks), SWE-bench (code generation), MATH (mathematical reasoning), and the Agent Security Benchmark (security). We translate each dataset into eleven diverse languages, resulting in 805 unique tasks and 9,660 total language-specific instances - enabling a systematic analysis of the multilingual effect on AI agents’ performance and robustness. Empirically, we observe degradation in both performance and security when transitioning from English to other languages, with severity varying by task and correlating with the amount of translated input. Building on these findings, we provide actionable recommendations to guide agentic AI systems development and assessment under multilingual settings. This work establishes the first standardized evaluation framework for multilingual agentic AI, encouraging future research towards equitable, reliable, and accessible agentic AI. MAPS benchmark suite will become publicly available upon publication.  [^1] https://huggingface.co/datasets/Fujitsu-FRE/MAPS
Abstract
This supplementary material provides additional details supporting our MAPS benchmark, including extended benchmark comparisons, experimental protocols, and multilingual evaluation results. It is organized as follows:
  • Reproducibility Resources – links to the full datasets and codebase, along with licensing terms (MIT) to enable open use and distribution for research and development purposes.
  • Agentic AI Benchmark Comparison – an expanded analysis of existing performance and security agentic benchmarks across dimensions such as evaluation objective, task type, answer format, and multilingual support.
  • Hybrid Translation Pipeline – additional information on our hybrid translation methodology, including domain-specific prompt design and translation configuration experiments.
  • Manual Evaluation of Translation Quality – annotation guidelines, evaluation examples, and a complete breakdown of human verification scores per dataset and language.
  • Experimental Settings – full specification of the technical setup, agent configurations, and language adaptation process used to isolate multilingual effects across tasks.
  • Extended Results – detailed multilingual performance tables for GAIA, MATH, SWE-Bench, and ASB across individual evaluation runs.
MAPS: A Multilingual Benchmark for Agent Performance and Security
Omer Hofman1, Oren Rachmil1, Shamik Bose1, Vikas Pahuja1, Jonathan Brokman1, Toshiya Shimizu2, Trisha Starostina3, Kelly Marchisio3, Seraphina Goldfarb-Tarrant3, Roman Vainshtein1 1Fujitsu Research of Europe  2Fujitsu Limited  3Cohere
Figure 1: MAPS benchmark suite evaluates LLM-based agents across $12$ languages and 44 agentic benchmarks covering performance and security.
1 Introduction
LLM-based agentic AI systems combine multi-step reasoning with external tools and memory to solve open-ended tasks such as code generation, web navigation, planning, and transactional services like booking and ordering (Acharya et al., 2025). By doing so, they extend to complex, real-world problems beyond standard LLM benchmarks. Since such real-world applications serve speakers of diverse languages, maintaining consistent reliability in every language becomes critical. However, since agentic behavior is grounded in LLMs, which often perform inconsistently across languages (Deng et al., 2023; Wang et al., 2023), agents may inherit these multilingual limitations as well, affecting their functionality and trustworthiness. This presents a barrier to equitable access, as non-English users may face degraded responses, incorrect tool actions, or unsafe behaviors—failures that can lead to actual harm in the real world, including erroneous transactions, data corruption, and security vulnerabilities (Zhang et al., 2024).
To assess emerging agentic systems, various benchmarks have been proposed to evaluate agent performance across a range of tasks (Mialon et al., 2023; Jimenez et al., 2023; Chang et al., 2024; Xu et al., 2024). However, these benchmarks remain English-only. In contrast to multilingual LLM benchmarks (Dang et al., 2024; Shi et al., 2022; Goyal et al., 2022), no equivalent exists for agentic AI—creating a blind spot in our understanding of cross-language performance, safety, and security.
In this paper, we address this gap. We hypothesize that multilingual settings will reveal performance and security gaps in agentic systems that are not captured by the existing, English-only benchmarks. To investigate this, we introduce MAPS, a Multilingual Agentic AI Benchmark Suite for Performance and Security. MAPS is based on four established agentic benchmarks: GAIA (real-world tasks) (Mialon et al., 2023), SWE-bench (code generation) (Jimenez et al., 2023), MATH (mathematical reasoning) (Hendrycks et al., 2021), and the Agent Security Benchmark (security) (Zhang et al., 2024). These benchmarks are extended to eleven typologically diverse languages beyond English [^2] MAPS offers evaluation in twelve languages: English, German, Spanish, Portuguese (Brazil), Japanese, Russian, Italian, Arabic, Hebrew, Korean, Chinese and Hindi., by employing a hybrid machine and LLM-based translation approach Ki and Carpuat (2024) with extended verification and enhancements. In total, MAPS includes 805 unique tasks, each available in 12 language versions—including the original English and 11 translated variants—for a total of 9,660 instances. An overview of the benchmark layout is shown in Fig. 1.
To demonstrate the use of MAPS and test our hypothesis, we selected a leading open-source agent associated with each of the four original benchmarks and applied it to the corresponding multilingual extension. We observed notable declines in both task performance and security when moving from English to other languages, with the severity of these drops varying by task type and correlating with the proportion of non-English input tokens, suggesting that multilingual performance interventions should be targeted based on input composition and task sensitivity. Beyond overall degradation, our findings reveal that multilingual inputs can amplify agentic vulnerabilities in safety-critical tasks, highlighting the need for multilingual risk assessment. These results empirically support our hypothesis and demonstrate the utility of MAPS for multilingual evaluation of agentic AI systems.
This paper makes three primary contributions:
  • To the best of our knowledge, we introduce the first multilingual benchmark suite for agentic AI, extending four widely used benchmarks into eleven typologically diverse languages for systematic performance and security assessment.
  • The efficacy and quality of the proposed benchmark are demonstrated through a large-scale evaluation of leading agents as well as human expert verification.
  • We present the first quantifiable analysis and evidence that multilingual settings reveal critical performance, safety, and security gaps in agentic systems, along with actionable recommendations for improving their development.
Figure 2: Overview of our multi-stage translation pipeline for agentic benchmark construction. We start with machine translation for structural alignment, followed by LLM-based verification and enhancement. This approach is adapted from Ki and Carpuat (2024) but extended with task-specific prompting and fallback mechanisms tailored to the requirements of agentic AI evaluation.
2 Background and Related Work
2.1 Agentic AI Benchmarks
With the rapid advancement of LLM-based agents, a diverse suite of benchmarks has been developed to assess their autonomy, tool use, planning, and memory integration (Yao et al., 2024; Xu et al., 2024; Yehudai et al., 2025). We organized these suites along three primary dimensions.
Evaluation objective: Performance-oriented benchmarks measure task completion, multi-step reasoning, and correct tool use (e.g., AgentBoard (Chang et al., 2024)), whereas security-focused suites probe robustness to adversarial inputs, jailbreaks, and unsafe behaviors (e.g., AgentHarm (Andriushchenko et al., 2024)). Agentic task scope: Full-agentic evaluations present only problem statements and expected outcomes, requiring end-to-end planning and execution (e.g., GAIA (Mialon et al., 2023)), while semi-agentic frameworks supply scaffolding, such as code templates or mock APIs, to isolate the LLM’s reasoning and tool-selection core (e.g., AppWorld (Trivedi et al., 2024). Design and evaluation characteristics: Most benchmarks span a limited set of domains (three to five use cases), typically including real-world information retrieval and navigation (e.g., AssistantBench (Yoran et al., 2024)), code generation (e.g., SWE-Bench (Jimenez et al., 2023)), reasoning and planning (e.g., MATH (Hendrycks et al., 2021), and security scenarios (e.g., Agent Security Benchmark (Zhang et al., 2024)). They enable objective measurement despite agents’ open-ended capabilities, by formulating tasks with definitive ground truth, allowing clear determination of success (Jimenez et al., 2023; Mialon et al., 2023). Further comparisons are provided in the Appendix.
While multilingual LLM benchmarks such as XTREME (Hu et al., 2020), FLORES (Goyal et al., 2022), and SIB-200 (Adelani et al., 2023) have enabled broad cross-lingual evaluation, they do not assess interactive decision-making, tool use, or task execution - the core elements of agentic systems. Thus, existing multilingual benchmarks fail to capture the complexities and vulnerabilities agents face in non-English settings - underscoring the need for multilingual agentic benchmarks.
2.2 LLMs Multilingual Limitations
Studies show that pre-trained LLMs struggle with non-English inputs, particularly in languages that are low-resource or typologically distant from English. Multilingual benchmarks (Hu et al., 2020; Liang et al., 2020) report accuracy drops when moving from English to languages such as Swahili or Nepali. These gaps stem from imbalanced training data, with $90\%$ English data, as well as challenges in tokenizing morphologically rich languages and limited multilingual fine-tuning data (Jha, 2024). Notably, even large models (e.g., GPT-4, Llama 405B) face a “cross-lingual knowledge barrier” (Hendrycks et al., 2020), (Chua et al., 2024).
LLMs also face security risks in multilingual contexts. Since most security efforts have been English-centric, models are more prone to generate policy-violating outputs when processing non-English prompts (Deng et al., 2023; Wang et al., 2023; Aakanksha et al., 2024). Multilingual safety alignment, have been shown to be effective but it also degrades general performance latency (Aakanksha et al., 2024).
Since agentic AI systems are based on these LLMs, we hypothesize that they inherit the same language-dependent performance and security limitations. These agents carry out real-world tasks , thus any inherited shortcomings can lead to severe consequences. To our knowledge, no systematic evaluation has examined these vulnerabilities in agentic systems. To address this gap, we introduce MAPS, a multilingual agentic benchmark suite.
3 MAPS: Multilingual Agentic AI Benchmark Suite
To support multilingual evaluation of agentic systems, we construct a benchmark suite by extending established English-language datasets into multiple languages. This process requires careful dataset selection, translation procedures that preserve semantic and structural integrity, and mechanisms for ensuring evaluation consistency. The following subsections detail our methodology for translation, benchmark construction, and dataset composition.
3.1 Translation Pipeline
Our translation pipeline combines automated translation with human expert verification - emphasizing both semantic and structural integrity. We combine Neural Machine Translation (NMT) which excels at preserving format and structure (Koehn and Knowles, 2017; Aharoni et al., 2019), and LLM translation for broader high-level capabilities. This extends Ki and Carpuat (2024), adding automated quality checks, fallbacks and human expert verification. The overall process is illustrated in Fig. 2. Below is a short summary - further details are in the Appendix.
Automated Translation. This system first computes the NMT translation, then improves it via LLM. let $T:\mathcal{S}\!\times\!\mathcal{L}\rightarrow\mathcal{T}$ map an English instruction $s\in\mathcal{S}$ and a target language $L_{t}\in\mathcal{L}$ to the translation $t\in\mathcal{T}$. We rely on five primitives: $\hat{t}=M(s,L_{t})$: neural MT system. $A(s,\hat{t},L_{t})\!\in\!\{0,1\}$: adequacy check of $\hat{t}$. $\Phi_{\text{direct}}(s,L_{t})$: direct LLM translation. $\hat{t}_{enh}=\Phi_{\text{enh}}(s,\hat{t},L_{t})$: LLM enhancement of $\hat{t}$. $I(s,\hat{t})\!\in\!\{0,1\}$: integrity check (hallucination, drift). The translation favors $\hat{t}_{enh}$ with fallbacks as follows:
Input: source text $s$, target language $L_{t}$
Output: translation $t$
1.  $\hat{t}\leftarrow M(s,L_{t})$
2.  if $A(s,\hat{t},L_{t})=0$ then return $\Phi_{\text{direct}}(s,L_{t})$ fallback 1
3.  $\hat{t}_{enh}\leftarrow\Phi_{\text{enh}}(s,\hat{t},L_{t})$
4.  if $I(s,\hat{t}_{enh})=0$ then return $\hat{t}$ fallback 2
5.  return $\hat{t}_{enh}$
Human Expert Verification. To ensure this pipeline reliability across languages and task types, we conducted human verification on a representative subset of translations. Evaluation setup and results are in Subsection 3.3.
3.2 Dataset Selection and Composition
Dataset Selection. To enable robust multilingual evaluation of agentic capabilities, we construct MAPS benchmark suite from established agentic AI benchmarks. Those were selected based on four criteria: (1) strong adoption within the research community, including prior use in agentic evaluation; (2) clearly defined, closed-form answers to enable controlled evaluation; (3) sufficient difficulty to challenge open-source agents without saturating performance; and (4) practical solvability, ensuring that multilingual degradation can be measured. Accordingly, we selected four datasets spanning real-world reasoning, software engineering, mathematical problem solving, and security assessment.
GAIA. GAIA (Mialon et al., 2023) is a benchmark designed to evaluate agents’ performance on real-world assistant tasks. It includes curated questions that require multi-step reasoning and autonomous use of tools such as web browsers, code interpreters, or document analyzers. Each question has a single correct answer, and responses are evaluated by an exact match to a reference output.
SWE-bench. SWE-bench (Jimenez et al., 2023) is a software engineering benchmark constructed from real GitHub issues and associated pull requests across popular Python repositories. Each task presents a bug report and a codebase snapshot, and requires the agent to evaluate whether a proposed patch correctly resolves the issue. We adopt the verified subset OpenAI (2024a), in which agents are tasked with validating a patch rather than generating one.
MATH. The MATH dataset (Hendrycks et al., 2021) includes high-school level mathematical problems across seven topics, including algebra, geometry, and calculus. Tasks are structured to require symbolic manipulation and multi-step reasoning. Agent responses are evaluated by exact match against a reference solution.
Agent Security Benchmark (ASB). ASB benchmark (Zhang et al., 2024) provides a structured evaluation of agent robustness against adversarial threats, including prompt injections, memory poisoning, and tool misuse. Agents interact with injected prompts or environments, and evaluation is based on whether safety policies are violated, measured by attack success rate and refusal rate.
Data Composition. The metadata below summarizes the multilingual extension, including language coverage, scale, and pre-processing.
Translated Languages. We selected the following eleven typologically and geographically diverse languages: German, Spanish, Portuguese (Brazil), Japanese, Russian, Chinese, Italian, Arabic, Hebrew, Korean, and Hindi. This selection enables the evaluation of agent performance across a wide range of scripts, linguistic structures, and regional user populations. Details on the specific dialects translated for each language are provided in the Appendix.
Dataset Handling. To preserve the integrity and utility of the original datasets, we applied only minimal and targeted modifications. Across all datasets, we appended translations without modifying or removing any original metadata (such as task type, difficulty level, tools available, etc). Domain-specific syntax—such as equations in MATH, code snippets in SWE-bench, and adversarial prompts in ASB—was preserved exactly, maintaining the original task structure and technical fidelity. For MATH and SWE-bench, which were not originally designed for agentic evaluation, we further applied selective filtering to retain only the most challenging tasks based on the task difficulty field. This follows common practice in prior work to align non-agentic datasets with agentic evaluation settings (Wu et al., 2023), ensuring meaningful agent evaluation while avoiding trivial cases.
Data Volume. To balance performance and security evaluation, our benchmark comprises $805$ tasks: $405$ from performance-oriented datasets (GAIA, SWE-bench, MATH) and $400$ from the Agent Security Benchmark. We selected $165$ tasks from GAIA (full validation set), $140$ high-difficulty tasks from MATH ($20$ per topic across $7$ topics), and $100$ hard and medium tasks from SWE-bench. The remaining $400$ tasks include all security-relevant prompts from ASB. Each task was translated into $11$ languages, and combined with the original English version, this results in a total of $9,660$ multilingual tasks across $12$ languages.
To validate the benchmark’s utility and examine multilingual effects, we applied a leading agent to each dataset. Full details and performance results are reported in Section 4.
Figure 3: Performance of open-source agents across languages on four agentic benchmarks: GAIA, SWE-bench, MATH, and ASB. Each bar represents the agent’s accuracy (or attack success rate in ASB) for a given language, with English shown as the baseline. Error bars indicate std across three runs. Performance differences reflect each agent’s degradation or resilience in multilingual settings.
3.3 Translation Implementation and Verification
Translation Implementation Details. For NMT we used Google Cloud (2024). For LLM translation we used Cohere’s command A Cohere et al. (2025) and gpt-4o with a designated system prompt. Structured expressions such as Code, URL and equations were identified via the provided special tokens provided by the datasets Zhang et al. (2024); Hendrycks et al. (2021), and masked accordingly in the translation process.
Quality Assurance. To assess translation quality, we implemented a verification protocol involving native-speaking experts. Each sample was rated by a bilingual annotator fluent in English and the target language on a $1\textendash 5$ Likert scale across three established criteria: adequacy (semantic fidelity), fluency (grammatical and stylistic naturalness), and formatting accuracy (preservation of LaTeX, code, etc.) Freitag et al. (2021). A fourth metric, answerability, measured whether the translation preserved intent well enough for the annotator to confidently answer the question as if it were in English. Annotator instructions are provided in the Appendix. To validate verification reliability, we embedded “honeypot” samples with intentional errors; annotators reliably flagged these cases, confirming attentiveness and quality control. In total, 2,200 translations (25% of the benchmark) were verified, proportionally sampled across datasets and languages - for comparability, see Xu et al. (2025), where 1,600/6,000 tasks where verified.
Evaluation results confirm high translation quality across the benchmark, with an answerability rate of 94.294.2%, and mean scores of 4.43 (adequacy), 4.59 (fluency), and 4.75 (formatting) on a 1–5 Likert scale, supporting the benchmark’s preservation of semantic fidelity, linguistic naturalness, and structural integrity. Because a sample of n=2,000n=2{,}000 gives a worst-case 95% margin of error of $\pm 2.2$ percentage points (derivation in Appendix), the true corpus-level answerability lies in $92.0,\,96.4]\%$, indicating robust translation quality even after accounting for sampling uncertainty. Full per-language results and analysis are included in the Appendix. To support high-precision use cases, we also release a “verified” subset, consisting of 190~190 translations per language that passed native expert review across all four datasets.
4 Experiments
4.1 Experimental Settings
Agent Assignment per Dataset. To demonstrate the utility of MAPS, we evaluate open-source agents on each multilingual dataset. We assigned a separate agent to each dataset, selected based on task relevance. While a unified agent would offer a more controlled evaluation, current systems lack the generalization needed to perform well across diverse tasks (Gioacchini et al., 2024; Chang et al., 2024). We employ each agent in its original configuration of prompts, system settings, tools and LLM backbone - all of which are multilingual models.
Figure 4: a) Multilingual effect as a function of the proportion of translated language tokens in input prompts. Each point represents a benchmark-agent pair, with the multilingual effect computed as the average relative degradation in performance or security across non-English languages. The trend suggests a correlation between input translation extent and multilingual vulnerability. b) Two snippets exemplify a low-translation prompt (MATH) and a high-translation prompt (GAIA), clarifying the x-axis percentages in panel (a) and showing how the proportion of natural-language tokens, rather than task difficulty alone, drives the observed multilingual effect.
For GAIA, we used the OpenDeepResearch agent (von Platen et al., 2024), which integrates retrieval, web browsing, and tool use to support real-world reasoning. For MATH, we adopted MathChat (Wu et al., 2024), a zero-shot agent combining multi-turn reasoning with Python execution and the Wolfram Alpha tool. For SWE-bench, we applied SWE-agent (Yang et al., 2024b), which enables autonomous software reasoning through repository navigation, file editing, and test execution. For ASB, we built on the original benchmark infrastructure (Zhang et al., 2024), eliciting synthetic adversarial tool-use in various agentic scenarios (e.g. travel-agent, legal consultant, system admin).
We use OpenDeepResearch with GPT o1, MATHChat with GPT-4, SWE-agent with GPT-4.1, and ASB with Qwen2. Full configuration detailsare provided in the Appendix.
Experiment Protocol. For each benchmark, the agent was evaluated three times in each of the $12$ target languages, yielding a total of $36$ runs per dataset. Fig. 3 reports the means and standard deviations over these runs.
Metrics. We adopt the original evaluation metrics from each benchmark to ensure consistency with prior agent evaluations. For MathChat (Math) and OpenDeepResearch (GAIA), we report the percentage of correct answers - with the latter matching either the English or translated reference. For SWE-agent, we report the percentage the percentage of submitted patches that successfully resolve the coding issue. For the ASB agent, we report the attack success rate (ASR). Additionally, we introduce a new metric: Multilingual Effect, which quantifies the performance or security gap between English and the average of all other languages. Given an evaluation metric $M$, the Multilingual effect is defined as follows:
$\text{Multilingual Effect }=\frac{1}{n}\sum_{i=1}^{n}M_{\text{lang}_{i}}-M_{\text{en}}$ (1)
Where $M_{en}$ denotes the performance in English, n is the number of non-English languages in the dataset (in our case $n=11$), and $M_{\text{lang}_{i}}$ represents the performance in the i-th non-English language.
4.2 Results
Fig. 3 presents the performance of open-source agents across all four benchmarks in English and the eleven target languages. In GAIA and ASB, we observe clear performance and security drops: non-English languages consistently underperformed compared to English, with reductions of up to 16% (GAIA) and a rise in vulnerability by up to 17% in ASB. Notably, SWE-bench and MATH exhibit only minor variation across languages, with most scores clustering around the English baseline. The Appendix extends this analysis with language-wise breakdowns of relative performance degradation.
These results reveal important differences in how multilingual degradation manifests across task types. Although all tasks require complex reasoning, some are more constrained than others. For instance, SWE-bench and MATH primarily involve structured code or mathematical expressions, placing less emphasis on natural language and more on Pythonic syntax or formal notation, thereby reducing the multilinguality effect. On the other hand, GAIA focuses on solving real-world tasks where understanding flexible natural language prompts is essential. Thus, the importance of the natural language problem statement is significantly higher. To explore this variation, we examine the proportion of localized, target-language text in each benchmark’s input (Fig. 4).
Interestingly, we observe that Chinese and Japanese yields the lowest ASR in the ASB benchmark, indicating the highest robustness to adversarial prompts in these two Asian languages. We attribute this pattern to the choice of backbone: the ASB agent was implemented using the Qwen2 model  (Yang et al., 2024a), which is known for its strong alignment for Chinese and Japanese language tasks. Qwen2 has consistently demonstrated strong performance in Chinese and Japanese-specific LLM benchmarks  [^3] https://rinnakk.github.io/research/benchmarks/lm/index.html, suggesting that alignment to a particular language and strategic backbone selection can significantly enhance resilience against multilingual adversarial prompts.
Fig. 4 examines the relationship between prompt composition and multilingual performance. part (a) shows a correlation between the percentage of non-English tokens in the input and the average performance gap (relative to English) across all four datasets. Benchmarks with higher proportions of localized, target-language-oriented input, such as GAIA and ASB, exhibit greater degradation, whereas SWE-bench, with predominantly English input (e.g., code), shows higher preservation. Part (b) demonstrate the differences between a MATH item with minimal translated text versus a GAIA query that is almost entirely translated.
English Mean (Non-English)
Orig. Trans.
GAIA (ACC) $\uparrow$ 47.4% 35.8% 41.3%
ASB (ASR) $\downarrow$ 48.8% 58.4% 56.6%
Table 1: Self-translation on GAIA and ASB. Making the agent aware of the multilingual context by self-translation of each non-English prompt recovers much of the lost performance/robustness, yet still leaves a noticeable gap to the native English baseline.
Self-translate Ablation. To further examine agents in multilingual settings, we followed the self-translate strategy of Etxaniz et al. (2024): each agent received one additional system message instructing it to translate any non-English prompt into English before solving the task, see results in Table 1. This led to improvements over direct multilingual inference, yet a clear gap to the English baseline persists, which may stem from the additional translation step introducing context loss or subtle distortions.
Figure 5: Cross-lingual planning consistency for OpenDeepResearch and SWE-Agent. For each agent, we identify tasks where the model succeeds in English but fails in all other languages, then compute the semantic similarity between the English instruction and the initial planning step in each language using multilingual embeddings. OpenDeepResearch exhibits a strong cross-lingual alignment gap: in 85.7% of cases, the English planning step is more semantically faithful to the instruction than its counterparts in other languages. In contrast, SWE-Agent is more multilingual robust, with English outperforming other languages in only 25% of cases.
Cross-Lingual Failure Analysis: To investigate why some agents exhibit a more pronounced multilingual effect than others, we compare their initial task planning across languages—specifically, the agent’s first planning step, where it breaks down the instruction into a sequence of actions to solve the task. We compute the cosine similarity between the English instruction and each language’s corresponding planning step using multilingual embeddings from OpenAI’s text-embedding-3-small model OpenAI (2024b). This allows us to quantify how closely the agent’s reasoning in each language aligns with the original task intent expressed in English.
Our analysis focuses on cases where agents succeed in English but fail in all other languages. We compare two representative agents: OpenDeepResearchAgent on the GAIA dataset and SWE-Agent on the SWE-Bench dataset, as they exhibit the strongest and weakest multilingual effects, respectively (see Fig. 4).
As shown in Fig. 5, OpenDeepResearchAgent displays a strong English bias in planning alignment on GAIA: planning steps in English are more semantically aligned with the instruction than in other languages (0.76 vs. 0.70 on average), with lower variance. This indicates that the agent’s reasoning is more faithful to the instruction in English and degrades across languages. In contrast, SWE-Agent exhibits much less cross-lingual variation on SWE-Bench: planning similarity remains consistent across languages, suggesting that its reasoning process is less sensitive to linguistic differences.
These findings suggest that when an agent’s initial planning remains stable across languages, the impact of the multilingual effect is reduced. For instance, SWE-Agent follows a structured and language-invariant approach—reproducing and executing code—which helps maintain alignment regardless of language. Conversely, GAIA relies more on open-ended natural language reasoning, making it more susceptible to semantic drift across translations.
5 Discussion
5.1 Guidelines for Multilingual Agent Deployment and Risk Assessment
Language-Composition-Aware Deployment Guidelines. Before deploying an AI agent in a multilingual setting, analyze the linguistic composition of its expected input, particularly the balance between structured elements (e.g., code, formal queries) and localized natural language. Inputs with a high proportion of natural language, tend to increase the risk of performance and safety degradation. We therefore recommend that in such case, developers should conduct a Multilingual Benchmark Assessment using a diverse, language-sensitive evaluation suite, such as ours, for AI agents operating across languages. This helps reveal vulnerabilities and promotes reliable real-world behavior across languages.
Multilingual Inputs Amplify Agentic Security Vulnerabilities. Our evaluation on ASB revealed that multilingual adversarial inputs can bypass agent safety mechanisms with minimal sophistication. Direct translations of English jailbreak prompts—without any adaptation or obfuscation—were sufficient to induce policy-violating behavior in multiple languages. This highlights a critical risk: localized inputs make simple attacks more effective, exposing security flaws in AI agents. Developers of safety-critical agentic systems should treat multilingual robustness as a core security concern and include translated prompts in safety evaluations using benchmarks like ours.
5.2 Benchmark Limitation
MAPS includes four datasets and eleven target languages, offering a strong foundation for multilingual evaluation. Nevertheless, extending these to additional domains, e.g. healthcare or legal reasoning, as well as to extremely low-resource languages (e.g., Amharic or Uyghur), would further enhance the benchmark. Nonetheless, the current suite already surfaces clear gaps across languages, offering valuable insights for multilingual deployment.
6 Conclusions
We introduce MAPS, the first multilingual benchmark suite for evaluating agentic AI systems, addressing a critical gap in assessing language-specific performance and safety limitations. By extending four widely used agentic benchmarks—GAIA, SWE-bench, MATH, and ASB—into eleven diverse languages, our suite enables the analysis of agent behavior under multilingual conditions. Constructed through a hybrid translation pipeline and human verification, MAPS ensures high linguistic and structural consistency. Experiments reveal critical gaps in performance and robustness when agents operate in non-English settings, particularly in tasks involving natural language reasoning and safety-critical behavior. These findings underscore the importance of language-aware evaluation and multilingual adaptation, especially for real-world agentic deployments. We view MAPS as a practical foundation for building more inclusive, resilient, and globally reliable agentic AI systems, and invite the community to extend it.
Appendix A Appendix
A.1 Reproducibility Resources
We release the full MAPS dataset, which includes translations of four agentic AI benchmarks (GAIA, SWE-bench, Math, and Agent Safety Benchmark) across 12 languages. A verified subset of MAPS, manually reviewed by native speakers, is also included for higher-quality evaluations. Links to both the full dataset and verified subset can be found at: MAPS Benchmark. We also provide the code for our hybrid translation pipeline, including prompt design and format-preserving techniques, available at: Hybrid Translation Implementation. All resources are released under the MIT License, allowing unrestricted use, modification, and distribution for academic and commercial purposes, provided proper attribution is given.
A.2 Agentic AI Benchmark Comparison
To contextualize MAPS within the growing landscape of agentic evaluation, we provide a detailed comparison of recent benchmarks across core design dimensions. While the main paper (Section 2.1) outlines a conceptual taxonomy—covering evaluation objectives, agent autonomy levels, and domain specialization—this supplementary section complements it with a structured, side-by-side table. Our comparison highlights key properties such as the ground truth availability, domain coverage, and multilingual support. Notably, MAPS remains the only benchmark to offer multilingual assessment across multiple agentic tasks, spanning both performance and safety. This contrast underscores an existing gap in multilingual agent evaluation and illustrates how MAPS fills it.
Paper Evaluation Objective Task Scope Domain Coverage # of tasks Answer Format GT Available Multilingual GAIA (Mialon et al., 2023) Performance Full-agentic General, multimodal 460 Close SWE Bench (Jimenez et al., 2023) Performance Full-agentic Software engineering 3000 Close MATH (Hendrycks et al., 2021) Performance Full-agentic Math solver 12,500 Close $\tau$-bench Yao et al. (2024) Performance Semi-agentic Airline, retail 165 Close AgentBoard (Chang et al., 2024) Performance Semi-agentic General – 9 varied tasks 134 Close App World (Trivedi et al., 2024) Performance Semi-agentic Apps, code execution, APIs 750 Close Partial TheAgentCompany (Xu et al., 2024) Performance Semi-agentic SW Dev, HR, Admin, DS 175 Close ARA (Kinniment et al., 2023) Performance, Security Semi-agentic SW Dev, phishing, web searches 12 Open TravelPlanner (Xie et al., 2024) Performance Full-agentic Travel planner, tourist 1225 Open
Table 2: Comparison of existing agentic AI benchmarks along evaluation objective, scope, domain coverage, and multilingual availability.
Paper Evaluation
Objective
Task Scope Domain Coverage # of tasks Answer Format GT Available Multilingual
Agent Security Bench (Zhang et al., 2024) Security Semi-agentic Offensive cyber operations, unsafe tool 90,000 Close
AgentHarm (Andriushchenko et al., 2024) Security Full-agentic Fraud, cybercrime, harassment 110 Open
Agent-SafetyBench (Zhang et al., 2024) Security Semi-agentic Jailbreaks, data leakage, unsafe tool use 2,200 Close
AgentDojo (Debenedetti et al., 2024) Security Semi-agentic Prompt injection, phishing, data leakage 97 Both Partial
InjecAgent (Zhan et al., 2024) Security Semi-agentic Indirect prompt injection, data leakage 1,000 Close
AgentPoison (Chen et al., 2024) Security Full-agentic Backdoor attacks, data poisoning ~20 Open
BIPIA (Yi et al., 2023) Security Semi-agentic Indirect prompt injection, data leakage 5 Close
Evil Geniuses (Tian et al., 2023) Security Full-agentic Prompt injection, data leakage 156 Open
Table 3: Comparison of agentic security benchmarks by scope, threat type, answer format, and multilingual coverage.
Table 2 surveys general-purpose agentic benchmarks, revealing that most focus exclusively on performance, operate within semi-agentic environments, and lack multilingual coverage. Full-agentic settings—those requiring autonomous planning and execution without scaffolding—are limited to a few benchmarks such as GAIA, SWE-Bench, and MATH, each addressing distinct domains: multimodal reasoning, software engineering, and mathematical problem solving, respectively. Table 3 focuses on security-oriented agentic benchmarks. While these benchmarks advance robustness evaluation, most adopt semi-agentic setups and open-ended answer formats, and often withhold ground truth to prevent overfitting.
Based on this analysis, we selected four complementary benchmarks—GAIA, SWE-Bench, MATH, and Agent Security Bench (ASB)—that collectively span performance and security objectives, cover diverse domains including real-world reasoning, code, math, and cybersecurity, and offer closed-form outputs with accessible ground truth. Notably, three of the four operate in fully agentic settings, aligning with our emphasis on end-to-end autonomous decision-making. This configuration enables reliable, multilingual evaluation across both agent capabilities and vulnerabilities, providing a robust foundation for MAPS.
A.3 Hybrid Translation Pipeline Implementation Additional Information
Figure 6: System prompt used for math translation refinement. This prompt is part of the hybrid translation pipeline and instructs the model to improve machine-translated math questions while preserving all LaTeX expressions and symbolic structure. As part of our hybrid translation pipeline, we designed domain-specific prompts for each benchmark to ensure appropriate evaluation and structural fidelity.
Reliable multilingual evaluation of agentic AI systems hinges on translating task instructions with both semantic and structural cross-language fidelity. Neural MT excels at preserving format and structure, but struggles in low-resource or specialized domains (Koehn and Knowles, 2017; Aharoni et al., 2019). Translation via instructed LLMs offers broader high-level capabilities at the cost of occasional hallucinations and semantic drift (Hendy et al., 2023; Yan et al., 2024). To balance these trade-offs, hybrid pipelines were suggested by Ki and Carpuat (2024); Mohammed (2025), combining format-preserving MT with LLM-based refinement. For MAPS, we extend Ki and Carpuat (2024): First, Ki and Carpuat (2024) was not designed with our benchmarks in mind, thus significant per-benchmark prompting had to be done. Second, we added automated quality checks, fallbacks, and expert verification to ensure the cross-language fidelity needed for agentic benchmarks (Fig. 2).
Formally, let us express our translation pipeline as a function $T:\mathcal{S}\times\mathcal{L}\rightarrow\mathcal{T}$, where $s\in\mathcal{S}$ is a task-instruction instance in source-language (English), $L_{t}\in\mathcal{L}$ is the target language, and $t\in\mathcal{T}$ is the resulting translated output. The pipeline begins with machine translation (MT) to establish a structural foundation: Denote $M(s,L_{t})$, the MT function, implemented as a high-quality, off-the-shelf NMT system. Its output provides a structurally faithful baseline for subsequent steps.
Next, we apply a verification step using an LLM to assess whether the translation adequately preserves the source meaning. This is modeled as a binary function $A(s,M(s,L_{t}),L_{t})\rightarrow{0,1}$, where the LLM compares the original and translated texts to detect major semantic errors or omissions.
Based on verification outcomes, the pipeline follows one of two distinct paths. If $A=0$ (indicating MT failure), the pipeline employs direct LLM translation: Denote $\Phi_{\text{direct}}(s,L_{t})$ the output of an LLM prompted to directly translate $s$ to language $L_{t}$ (without using the MT output). If $A=1$ (indicating acceptable MT), an LLM enhances the translation while preserving its basic structure: Denote $\Phi_{\text{enhancement}}(s,M(s,L_{t}),L_{t})$ as the output of an LLM, guided to improve the MT output while maintaining structural consistency.
To ensure semantic integrity, we apply a second binary check: $I(s,\Phi_{\text{enhancement}})\rightarrow{0,1}$, targeting common LLM failure modes, such as hallucinations, omissions, misinterpretations (e.g., answering instead of translating), and semantic drift. If this verification fails, we revert to the original MT output (which passed the initial verification test).
These conditional steps form a robust decision framework: If MT is rejected, we use a direct LLM translation; if accepted but enhancement fails integrity verification, we fall back to the MT; Otherwise, we use the enhanced translation. Importantly, since both MT and LLM-based translation are robust technologies, we did not encounter any scenario in which both systems failed to produce a valid translation. A formal expression of this pipeline is provided in the Appendix.
For the translation pipeline, we used Python 3.11 alongside a hybrid setup comprising Google’s Neural Machine Translation (NMT) API, Cohere Command-A, and GPT-4o. To support high-quality multilingual alignment, we crafted dedicated system prompts for each stage of the pipeline—translation verification, semantic enhancement, and structural integrity checks—tailored to the specific requirements of each domain. For example, GAIA tasks involve real-world decision-making, where preserving subtle intent and contextual detail across languages is critical. In contrast, MATH tasks require precise handling of symbolic notation and mathematical terminology, necessitating strict preservation of LaTeX expressions and domain-appropriate language. An example of a task-specific system prompt is provided in Fig. 6. The full translation pipeline implementation, including prompt templates and evaluation logic, is available  [^4] https://tinyurl.com/hybrid-translation-pipeline.
To evaluate translation quality across different configurations, we employed LLM-based evaluation using a structured prompt. This approach follows the framework introduced by Jiang et al.(Kocmi and Federmann, 2023), who show that LLMs can reliably assess translation quality when guided by clear criteria. We adapted and extended their methodology to include four dimensions tailored to our benchmark context (those metrics are similar to the ones we define in the manual verification experiment): (1) adequacy—semantic fidelity to the original task, (2) fluency—naturalness and grammaticality in the target language, (3) formatting accuracy—preservation of structural elements such as LaTeX, code, or tool references, and (4) answerability—whether the translation remains usable and complete. The evaluation prompt used is provided in Fig.7. Unlike traditional evaluation metrics such as COMET (Rei et al., 2020), which rely on the same-language reference translations and embedding-based comparisons, our agentic tasks often lack ground truth references and span cross-language instruction. In such settings, reference-based metrics become inapplicable, whereas LLM-based zero-shot evaluation enables context-sensitive, reference-free scoring.
Using the LLM as a judge approach, we evaluated four translation strategies (Table 4) by translating the GAIA validation set, to empirically identify the most effective pipeline design. This aligns with the analysis in Section 3 of the main manuscript, which discusses the trade-offs between machine translation (MT) and LLM-based generation. As expected, Google NMT showed strong performance on formatting but slightly lagged in semantic adequacy and fluency. Command-A improved fluency and meaning preservation but occasionally introduced structural drift. Hybrid approaches—where NMT outputs were refined by either GPT-4o or Command-A—consistently outperformed single-method setups. Notably, the Google+Command-A variant achieved the highest scores across all metrics, including adequacy (4.76), fluency (4.90), formatting accuracy (4.96), and answerability (95.9%). These results validate our decision to adopt a hybrid approach and extend prior work(Ki and Carpuat, 2024) by integrating domain-specific prompting, automated checks, and structure-aware refinement for multilingual translation in agentic AI benchmarks.
Translation Method Adequacy (1-5) Fluency (1-5) Formatting Accuracy (1-5) Answerable (%)
Google Translate NMT 4.667 4.870 4.946 92.9%
Command A (LLM only) 4.704 4.872 4.934 93.9%
Hybrid (Google Translate NMT + GPT-4o) 4.709 4.832 4.923 94.3%
Hybrid (Google Translate NMT + Command A) 4.760 4.899 4.959 95.9%
Table 4: Translation quality comparison across methods based on human evaluation. Bolded values represent the best-performing configuration for each metric.
Figure 7: System prompt used to guide the LLM-based translation evaluation. The prompt instructs the model to assess translations across four dimensions—adequacy, fluency, formatting accuracy, and answerability—and return results in a strict JSON format.
A.4 Hybrid Translation Manual Evaluation Additional Information
To ensure high-fidelity multilingual translation in MAPS, we conducted a human evaluation of translated samples using a standardized annotation protocol. This procedure mirrors the evaluation framework used for our LLM-based assessment, following the methodology proposed by Jiang et al. (Kocmi and Federmann, 2023). Below, we reproduce the instructions provided to annotators, which guided their evaluation of semantic adequacy, linguistic fluency, formatting consistency, and answerability across all benchmark domains.
A.4.1 Translation Verification Guidelines for Annotators
Purpose. You are asked to verify the quality of translations from English into your target language for a multilingual benchmark dataset. Your careful evaluation ensures that the translations are accurate, fluent, and correctly formatted.
Dataset Overview and Translation Challenges. You will review 200 translated samples drawn from four benchmark datasets:
  • GAIA (50 samples): natural language questions requiring web search and file analysis.
  • Agent Security Benchmark (60 samples): tasks probing unsafe or manipulative agent instructions.
  • MATH (70 samples): algebra, calculus, probability, and geometry problems, often containing LaTeX and variables.
  • SWE-Bench Verified (20 samples): GitHub issues involving bug descriptions and code snippets.
Evaluation Procedure. For each sample:
  1. 1.
    Read the English source text.
  2. 2.
    Read the translated text in the target language.
  3. 3.
    Assign the following ratings:
    • Adequacy (1–5): how well the meaning is preserved.
    • Fluency (1–5): how natural and grammatically correct the translation is.
    • Formatting Accuracy (1–5): whether LaTeX, variable names, units, and structure are preserved.
    • Answerable? (Yes/No): Answer No if you feel you cannot answer the original question because of a major translation mistake (e.g., reversed meaning, missing information). Otherwise, mark Yes.
    • Notes (optional): any remarks you wish to add.
Rating Scales.
  • Adequacy: 5 = perfect meaning preservation, 1 = major meaning loss.
  • Fluency: 5 = native-like fluency, 1 = broken or ungrammatical.
  • Formatting Accuracy: 5 = all formatting correct, 1 = major formatting errors.
Important Notes.
  • Do not correct or rewrite the translations.
  • Pay special attention to LaTeX symbols, variable names, and numerical units.
Sheet Columns Example (one row per sample):
  • Item ID – Sample identifier.
  • English Source Text – Original English input.
  • Translated Text – Translated version.
  • Adequacy (1–5) – Meaning preservation.
  • Fluency (1–5) – Naturalness and grammar.
  • Formatting Accuracy (1–5) – Structural fidelity.
  • Answerable? – Unanswerable due to error?
  • Notes (Optional) – Free-text comments.
Figure 8: Translation verification examples submitted by annotators during manual evaluation. Left: A successful Japanese translation from the GAIA dataset, rated highly across all dimensions with only a minor terminology note. Right: A Korean MATH example flagged as critically erroneous due to semantic and grammatical mismatches in imperative phrasing. These examples illustrate the evaluation schema’s ability to surface both subtle formatting issues and major adequacy failures.
A.4.2 Translation Verification Extended Results
To ensure the validity of our multilingual benchmark, human verifiers manually reviewed approximately 25% of the translated data—corresponding to over 2,000 samples—across all tasks and languages. In the main manuscript, we reported the overall scores for adequacy, fluency, formatting accuracy, and the error rate derived from answerability. In Tables 6, 7, 8 and 9 we present the full breakdown of results for each individual dataset and in Table 10 we conclude with the manual verification average results across all datasets. Across all four benchmark datasets, native verifiers’ ratings reveal high translation quality, with average adequacy (4.47), fluency (4.60), and formatting accuracy (4.76) on a 1–5 scale. Notably, answerability remained consistently high (mean: 0.94), suggesting that most translations retained functional fidelity. Formatting consistency was strongest for ASB and GAIA, while fluency slightly declined for Arabic and Japanese, correlating with minor adequacy drops. Additionally, Fig. 8 presents one successful and one failed translation example, as annotated by Japanese and Korean evaluators, respectively. These results confirm that closed-form multilingual alignment is achievable across diverse tasks without compromising validity.
A.4.3 Statistical sufficiency of the 25% manual audit
Let n=2,000n=2{,}000 be the size of our manually verified subset ($\approx 25\%$ of the full benchmark). For any binary outcome—here, Answerable vs. Unanswerable—the sample proportion $\hat{p}$ is an unbiased estimator of the true corpus proportion $p$. Under the binomial model, its standard error is
$\mathrm{SE}(\hat{p})=\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}.$
The variance is maximised when $\hat{p}=0.5$, giving a worst–case standard error $\mathrm{SE}_{\max}=0.5/\sqrt{n}$. Using the normal approximation with critical value $z_{0.025}=1.96$, the 95 % two–sided margin of error is therefore
$E=1.96\times\frac{0.5}{\sqrt{n}}=\frac{0.98}{\sqrt{n}}=\frac{0.98}{\sqrt{2{,}000}}\approx 0.022.$
Hence any proportion reported from the manual audit has a 95 % confidence interval of at most $\pm 2.2$ percentage points. Put differently, any systematic translation error affecting more than $\approx 2\%$ of the corpus would almost certainly have been detected in our audit. Because the sample was stratified by language and dataset, the same formula applies within each stratum with $n_{\ell}\!=\!0.25\,N_{\ell}$ items audited per language $\ell$. Putting the numbers together, the overall Answerable estimate $\hat{p}=0.944$ therefore has a 95 % confidence interval
$0.944\pm 0.022\;=\;[0.922,\;0.966].$
Thus, even in the worst case, the corpus-level answerability remains above 92 %.
Language Variant
Chinese Simplified (Mainland)
Russian Standard (Russia)
Arabic Modern Standard Arabic
Hebrew Modern Israeli
German Standard German
Spanish Latin American
Hindi Standard Hindi
Japanese Standard Japanese
Korean Standard Korean
Italian Standard Italian
Portuguese Brazilian Portuguese
Table 5: Dialect used per language.
GAIA
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.14 4.04 4.74 0.91
ja 4.30 4.70 4.91 0.88
he 4.20 4.62 4.86 0.88
ko 4.49 4.84 4.40 0.94
pt-br 3.93 4.16 4.53 0.96
it 4.26 4.24 4.72 0.91
es 4.18 4.24 4.99 0.94
ar 3.54 3.79 4.41 0.98
hi 4.22 4.52 4.72 0.92
ru 4.84 4.99 4.90 0.96
AVG 4.21 4.41 4.72 0.93
Table 6: Manual translation evaluation scores for the GAIA dataset across 10 languages. Metrics include adequacy, fluency, formatting accuracy (all rated on a 1–5 scale), and answerability (proportion of translations judged as answerable).
MATH
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.80 4.87 4.97 1.00
ja 4.11 4.41 4.96 0.82
he 4.61 4.46 4.79 0.94
ko 4.62 4.95 4.66 0.93
pt-br 4.95 4.84 4.62 1.00
it 4.65 4.85 4.88 1.00
es 4.74 4.49 4.28 0.87
ar 3.78 3.99 3.81 0.82
hi 4.76 4.79 4.96 0.95
ru 4.75 5.00 4.75 0.99
AVG 4.58 4.67 4.67 0.93
Table 7: Manual translation evaluation scores for the SWE-Bench dataset across 10 languages.
SWE-Bench
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.30 4.80 4.50 1.00
ja 4.40 4.05 5.00 0.85
he 4.55 4.50 4.50 0.90
ko 4.35 4.25 4.40 0.90
pt-br 4.75 4.80 4.75 1.00
it 4.55 4.65 4.35 0.95
es 4.00 4.50 4.50 0.95
ar 3.60 4.30 4.20 0.90
hi 4.40 4.35 4.80 0.90
ru 4.85 4.95 4.80 0.95
AVG 4.375 4.515 4.580 0.930
Table 8: Manual translation evaluation scores for the SWE-Bench dataset across 10 languages.
ASB
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.10 4.53 4.97 0.95
ja 4.68 4.76 4.91 0.96
he 4.68 4.67 5.00 0.92
ko 4.95 4.93 4.96 0.98
pt-br 4.72 4.55 4.92 1.00
it 4.25 4.47 4.80 0.98
es 4.57 4.40 4.87 0.97
ar 4.45 4.68 4.90 0.98
hi 4.88 4.85 4.95 0.98
ru 4.82 4.90 4.83 0.98
AVG 4.61 4.67 4.91 0.97
Table 9: Manual translation evaluation scores for the ASB dataset across 10 languages.
All Benchmarks (Average)
Lang Adequacy (1–5) Fluency (1–5) Formatting Accuracy (1–5) Answerable
de 4.375 4.554 4.867 0.964
ja 4.356 4.553 4.936 0.881
he 4.525 4.565 4.840 0.915
ko 4.660 4.847 4.660 0.943
pt-br 4.609 4.580 4.703 0.989
it 4.422 4.565 4.763 0.968
es 4.475 4.403 4.657 0.922
ar 3.859 4.255 4.423 0.936
hi 4.625 4.697 4.881 0.947
ru 4.803 4.961 4.817 0.974
AVG 4.471 4.598 4.755 0.944
Table 10: Average manual translation evaluation scores across all benchmarks (GAIA, MATH, SWE-Bench, and ASB). Metrics include adequacy, fluency, and formatting accuracy (rated on a 1–5 scale), and answerability (proportion of samples deemed answerable).
A.4.4 Language Variants Used
To ensure wide applicability across regions, we translated the dataset into 11 languages using the most widely recognized and formal variant in each case. The table below lists the dialect or regional standard used in the translation process.
Figure 9: Trajectory variability between English and Hindi agents in solving the same question. The English agent follows a computation-driven path—extracting raw values and calculating the result—whereas the Hindi agent relies on surface-level textual matching and assumed equivalence. This leads to reasoning failure due to the absence of explicitly stated percentages in the Hindi query context.
Figure 10: Relative performance differences from English for each language, broken down by dataset. Negative values indicate a drop in performance compared to English, while positive values (notably in ASB) represent increased vulnerability. The trend highlights how multilingual effects vary by language and task type.
A.5 Experimental Settings Additional Information
We conducted all experiments on a Linux-based machine (Ubuntu 22.04) equipped with 64 virtual CPUs and 256 GB of RAM. To ensure reproducibility and compatibility, each agent was executed within its own isolated software environment tailored to its specific dependencies. To ensure a fair evaluation of multilingual effects (see Eq. 2 in the main manuscript), we preserved each agent’s original implementation without introducing any code modifications. This isolation strategy allowed us to assess the impact of language variation on agent performance and behavior without confounding factors from architectural or behavioral changes. Datasets were also left unchanged in structure, format, and content. Only the task inputs and expected outputs were translated into the target languages. This process yielded 12 language-specific evaluation suites per dataset (including English), each of which can be executed independently, enabling controlled multilingual evaluation with minimal setup and without requiring code adaptation.
A.5.1 Agents Configuration Additional Information
For the GAIA benchmark, we used the publicly available agent from OpenDeepResearch (von Platen et al., 2024), which features a modular framework supporting document retrieval, browser-based execution, and web search. The agent is designed for general-purpose, real-world reasoning and incorporates memory and planning via tool-augmented prompting. In our evaluation, we disabled the web search functionality due to its reliance on the external Serper API, which imposes rate limits and introduces latency variability. This omission was applied across all languages to ensure consistency, and throughput scalability under multilingual evaluation. For MATH, we adopted MathChat (Wu et al., 2024), a mathematical reasoning agent that combines multi-turn chain-of-thought prompting with symbolic computation using Python and the Wolfram Alpha API tool.
For SWE-bench, we used SWE-agent (Yang et al., 2024b), an autonomous software engineering agent capable of navigating repositories, modifying code, and executing test suites. It employs a plan-execute loop enhanced with file-level attention and dynamic tool invocation. For ASB, we built on the original benchmark infrastructure (Zhang et al., 2024), eliciting synthetic adversarial tool-use in various agentic scenarios (e.g. travel-agent, legal consultant, system admin).
Each of the above was executed using the same LLM backbone reported in its original publication. All selected models are considered to be multilingual models. Specifically, GAIA used GPT-O1 (model version: 2024-12-17), MATHChat used GPT-4 (model version: turbo-2024-04-09), SWE-agent used GPT-4.1 (model version: 2025-04-14), and ASB used Qwen2 7B (model version: 2024-10-01).
A.6 Results Additional Information
This section supplements the main manuscript with detailed quantitative and qualitative results. While the main paper summarizes overall trends and multilingual effects, Tables 1114 provide per-language scores, averages, and standard deviations for each benchmark. Fig. 9 provides a multilingual qualitative failure example of Hindi language in GAIA, including a comparison of main steps in English and Hindi. Fig.  10 provides a language-wise breakdown of relative performance degradation across datasets. From the Figure, we can see that there is no clear correlation between multilingual security robustness (ASB) and multilingual performance degradation. This disconnect is especially clear in real-world, language-heavy tasks like GAIA, where performance drops sharply, while structured tasks like SWE-bench and MATH remain largely unaffected. This highlights that multilingual security alignment does not directly track with multilingual task accuracy, notably in language-rich agentic tasks.
Language Run #1 Run #2 Run #3 Mean StdDev
English (baseline) 0.5094 0.4450 0.4695 0.4746 0.0325
Portuguese (br) 0.3902 0.3658 0.3597 0.3719 0.0161
German 0.3658 0.3475 0.3597 0.3577 0.0093
Spanish 0.3475 0.3536 0.3680 0.3564 0.0105
Italian 0.3536 0.3353 0.3170 0.3353 0.0183
Russian 0.4268 0.4085 0.3530 0.3961 0.0384
Arabic 0.2866 0.3292 0.3170 0.3109 0.0219
Hebrew 0.3818 0.3597 0.3353 0.3589 0.0233
Hindi 0.3353 0.3780 0.3353 0.3496 0.0247
Korean 0.3597 0.3231 0.3475 0.3434 0.0186
Japanese 0.4180 0.3597 0.3170 0.3649 0.0507
Table 11: GAIA benchmark performance across languages.
Language Run #1 Run #2 Run #3 Mean StdDev
English (baseline) 0.29 0.33 0.31 0.31 0.020
Portuguese (br) 0.32 0.29 0.25 0.323 0.0351
German 0.29 0.27 0.30 0.287 0.0153
Spanish 0.31 0.31 0.31 0.31 0.0000
Italian 0.30 0.29 0.32 0.303 0.0150
Russian 0.31 0.34 0.34 0.33 0.0173
Arabic 0.33 0.28 0.31 0.307 0.0252
Hebrew 0.31 0.27 0.27 0.283 0.0231
Hindi 0.28 0.28 0.28 0.280 0.0000
Korean 0.30 0.32 0.32 0.313 0.0115
Japanese 0.31 0.29 0.29 0.297 0.0153
Table 12: SWE-Bench benchmark performance across languages.
Language Run #1 Run #2 Run #3 Mean StdDev
English (baseline) 0.514 0.514 0.514 0.514 0.000
Portuguese (br) 0.514 0.514 0.514 0.514 0.000
German 0.500 0.500 0.500 0.500 0.000
Spanish 0.536 0.536 0.536 0.536 0.000
Italian 0.479 0.479 0.479 0.479 0.000
Russian 0.507 0.507 0.507 0.507 0.000
Arabic 0.514 0.514 0.514 0.514 0.000
Hebrew 0.521 0.521 0.521 0.521 0.000
Hindi 0.436 0.443 0.443 0.440 0.004
Korean 0.414 0.421 0.421 0.419 0.004
Japanese 0.443 0.443 0.443 0.443 0.000
Table 13: MATH benchmark performance across languages.
Language Run #1 Run #2 Run #3 Mean StdDev
English (baseline) 0.534 0.536 0.532 0.534 0.002
Portuguese (br) 0.648 0.657 0.654 0.653 0.005
German 0.659 0.652 0.654 0.655 0.004
Spanish 0.595 0.613 0.597 0.602 0.010
Italian 0.612 0.618 0.603 0.611 0.008
Russian 0.636 0.646 0.642 0.641 0.005
Arabic 0.515 0.504 0.511 0.510 0.005
Hebrew 0.558 0.571 0.572 0.567 0.008
Hindi 0.617 0.616 0.614 0.616 0.002
Korean 0.590 0.589 0.588 0.589 0.001
Japanese 0.484 0.489 0.497 0.490 0.007
Table 14: ASB benchmark performance across languages.